RC meeting and Run Plan, Thursday February 16th
From SBS wiki
Return to the RC Meeting Calendar
Contents
RC Report
- https://logbooks.jlab.org/entry/4131947
- RC: Bogdan Wojtsekhowski, bogdanw@jlab.org
Summary
- Ran with ~30 uA beam overnight increasing to 40 uA
- No beam since 7am, up to now tune was not productive, Yves is at MCC
Issues
- DAQ lost time reach 1+ hour during owl shift
Current Status
- Beam tune
Run Plan
- Continue running 45 uA on He3
- Accelerator will prioritize beam and not have any maintenance days this week
Data Collection and Polarization Status
- 33.2C of uncorrected charge for this kinematic (38.8C of corrected charge)
- Last NMR 53.9%%
Other Topics/Status Updates
- Anything else?
- Bill wants to change the raster settings after some analysis
Requested Accesses to the Hall
- Jack would like to collect some cables